XYZ Co. has forecasted June sales of 600 units and July sales of 1000 units. The company maintains ending inventory equal to 125% of next month's sales. June beginning inventory reflects this policy. What is June's required production? 

Select one: 

a. 1,100 units b. 0-units c. 500 units d. 400 units 
Answer #1

Question 1

Option A is correct, i.e. 1,100 units

Opening inventory of June = 600 * 125% = 750

Closing inventory of June = 1,000 * 125% = 1,250

Production of June = Sales for the month of June + Closing inventory of June - Opening inventory of June

= 600 + 1,250 - 750

= 1,100 units